# Learning Linear Utility Functions From Pairwise Comparison Queries

### v1

@@ 1 Introduction
We first investigate the “passive” Probably Approximately Correct (PAC) learning model in which pairs of inputs (outcomes) are generated according to an unknown distribution $\mathcal{P}$, and responses to these comparison queries may be corrupted by noise that comes from a given distribution $\mathcal{Q}$. We show that when there is no noise in query responses, we can indeed efficiently learn to predict responses to unseen pairwise comparison queries with respect to the true utility function in the PAC sense. However, the presence of noise with any c.d.f. continuous at 00 leads to an exponential lower bound on sample complexity in this model. Nevertheless, we show that if the the input distribution is well-behaved and the query noise satisfies the Tsybakov noise condition, linear utilities are efficiently learnable in this model.
@@ 2 Related Work
Learning Utility Models from Pairwise Comparisons: The fundamental framework that we build on is random utility model learning, in which utility information is provided indirectly through ranking (e.g., pairwise) comparisons [23, 8, 35]. A common approach for learning random utility models from pairwise comparisons is by using maximum likelihood estimation (MLE) [26, 36, 35]. Unlike our setting, these typically investigate problems with a finite set of alternatives [31, 24]. Our focus, in contrast, is to learn a utility function over a continuous vector space. This problem has been considered by [42], who show that linear utilities can be efficiently learned using maximum likelihood estimators in the Bradley-Terry noise model.
PAC Learning of Halfspaces With Noise: Common models of noise in random utility models have the property that the closer the pair’s difference is to the halfspace defined by the linear utility function, the higher the chance that this comparison is flipped. This kind of noise has been previously studied as boundary consistent noise in [15, 39, 38]. The possibility of an arbitrarily small difference between noise and $\frac{1}{2}$ makes learning very challenging, even for learning halfspaces [4]. Consequently, proposals for tractable learning often rely on assumptions regarding distributions of noise. The spectrum of such noise models starts from the easiest version called random classification noise (RCN) [2] – where each label is flipped independently with equal probability exactly $\eta<\frac{1}{2}$ – where efficient algorithms exist [6]. At the other end are malicious noise [19] and agnostic learning [20, 21, 11, 13], where an adversarial oracle can flip any small constant fraction of the labels, typically making these models less tractable. Additionally, active learning has been proposed as a solution to reduce the the number of samples the learner needs to query before approximately learning the concept class with high probability. While active learning requires exponentially fewer labeled samples than PAC-learning for simple classes such as thresholds in one dimension, it fails in general to provide asymptotic improvement for broad classes such as halfspaces [12].
@@ 3 Preliminaries
We further define the difference between a pair $(\phi(x),\phi(x^{\prime}))$ as $\Delta_{\phi}(x)\equiv(\phi(x^{\prime})-\phi(x))\in[-1,1]^{m}$. Since $\tilde{u}(x^{\prime})-\tilde{u}(x)=w^{T}\Delta_{\phi}(x)+\tilde{\zeta}(\phi(x^{\prime}))-\tilde{\zeta}(\phi(x))$, it will be most useful to consider $\zeta\equiv\tilde{\zeta}(\phi(x^{\prime}))-\tilde{\zeta}(\phi(x))$ for a pair of feature vectors $(\phi(x),\phi(x^{\prime}))$, so that $Z_{\tilde{u}}(x,x^{\prime})=\mathrm{sign}(w^{T}\Delta_{\phi}(x)+\zeta)$. Let $\mathcal{Q}$ be the distribution over $\zeta$ induced by $\tilde{\mathcal{Q}}$, and let $F(\zeta)$ be its c.d.f., which we assume is continuous on $\mathbb{R}$ except in the noise-free special case. Then the probability that $y=1$ for a pair $(\phi(x),\phi(x^{\prime}))$, that is, the probability that $x^{\prime}\succ x$, is $\Pr(x^{\prime}\succ x)=\Pr(\zeta\leq w^{T}\Delta_{\phi}(x))=F(w^{T}\Delta_{\phi}(x))$.
@@
Our learnability discussion will be based on an adaptation of the Probably Approximately Correct (PAC) Learning framework [33], with the goal of identifying a utility model $\hat{u}$ that has error at most $\epsilon$ with probability at least $1-\delta$ for $\epsilon>0$ and $\delta>0$. We now formalize this as PAC learnability from pairwise comparisons (PAC-PC). Let $\mathbb{N}$ be the space of natural numbers.
Definition 1 (PAC-PC).
Given a noise distribution $\mathcal{Q}$, a utility class $\mathcal{U}$ is PAC learnable from pairwise comparisons (PAC-PC learnable) for error function $e(\hat{u},u)$ if there is a learning algorithm $\mathcal{A}$ and a function $n_{A}:(0,1)^{2}\rightarrow\mathbb{N}$ such that for any input distribution $\mathcal{P}$, and $\forall\epsilon,\delta\in(0,1)$, whenever $n\geq n_{A}(\epsilon,\delta)$, $\mathcal{A}(\{(x_{i},x_{i}^{\prime},y_{i})\}_{i=1}^{n})$ with $(x_{i},x_{i}^{\prime})\sim\mathcal{P}$ i.i.d. and $y_{i}=\mathrm{sign}(u(x_{i}^{\prime})-u(x_{i})+\zeta_{i})$ for $\zeta_{i}\sim\mathcal{Q}$ i.i.d. returns $\hat{u}$ with probability at least $1-\delta$ such that $e(\hat{u},u)\leq\epsilon$. Moreover, if $n_{A}$ is polynomial in $1/\epsilon,1/\delta,\mathrm{Dim}(\mathcal{U})$, where $\mathrm{Dim}(\mathcal{U})$ is the dimension of $\mathcal{U}$, then we say $\mathcal{U}$ is efficiently PAC-PC learnable.
@@ 4.1 Predicting Pairwise Preferences
Suppose $\zeta=0$. Then linear utility functions are efficiently PAC-PC under the error function $e_{1}$.
@@
Let $y=g(\Delta_{\phi}(x))=\mathrm{sign}(w^{*^{T}}\Delta_{\phi}(x))$. Then PAC-PC learning under $e_{1}$ and $\mathcal{P}$ is equivalent to PAC learnability of halfspaces $g(\Delta_{\phi}(x))$ under $\mathcal{P}^{\prime}_{\phi}$. As shown by Blumer et al. [7], halfspaces are learnable with sample complexity $\mathcal{O}(\frac{1}{\epsilon}(m+\log(\frac{1}{\delta})))$. ∎
@@
Due to the continuity of $F$ at 00, for any $t>0$, there exists $s>0$ such that for all $|v-0|\leq s$, $|F(v)-F(0)|\leq t$. As $F$ is monotonically increasing, and $-|w^{*^{T}}\Delta_{\phi}(x)|<0$, $\eta(\Delta_{\phi}(x))<F(0)=\frac{1}{2}$. Hence for a fixed $\epsilon<1$, we can pick $t_{0}=1/\exp(1/\epsilon)$. Then there exists a corresponding $s_{0}$ such that for a $(\phi(x),\phi(x^{\prime}))$ with $\Delta_{\phi}x$ of norm at most $s_{0}$, for all $w$, $|w^{T}\Delta_{\phi}(x)-0|\leq s_{0}$, $|\eta(\Delta_{\phi}(x))-\frac{1}{2}|\leq t_{0}.$ Now consider the oracle who always chooses one such pair $(\phi(x),\phi(x^{\prime}))$. With fixed probability at most $\frac{1}{2}+t_{0}$ it outputs the true label and with probability $\frac{1}{2}-t_{0}$ it outputs the false label. The task of this algorithm becomes to decide whether or not this single point $\Delta_{\phi}(x)$ is below or above the unknown halfspace $w,$ which is the same as detecting which way a coin is biased. As is well known, this may be determined by Chernoff’s inequality for Bernoulli distributions, $\Pr(S_{n}<n/2)=\Pr(S_{n}<(1-\frac{2t_{0}}{1+2t_{0}})n(\frac{1}{2}+t_{0}))\leq\exp(-\frac{n(\frac{2t_{0}}{1+2t_{0}})^{2}(\frac{1}{2}+t_{0})}{2})<\delta$, giving sample complexity $\mathcal{O}(\frac{1}{t_{0}}\log(\frac{1}{\delta}))=\mathcal{O}(\exp(\frac{1}{\epsilon})\log(\frac{1}{\delta}))$, which is exponential in $\frac{1}{\epsilon}$.∎
@@
[14] Let $C$ be a concept class of Boolean-valued functions over $X=\mathbb{R}^{d}$, $F$ be a family of distributions on $X$, $10<\epsilon<$ 1 be the error parameter, and $10\leq\alpha<$ 1, $A>0$ be parameters of the noise model. Let $f$ be an unknown target function in $C$. A Tsybakov example oracle, $EX^{Tsyb}(f,F)$, works as follows: Each time $EX^{Tsyb}(f,F)$ is invoked, it returns a labeled example $(x,y)$, such that: (a) $x\sim\mathcal{D}_{x}$, where $\mathcal{D}_{x}$ is a fixed distribution in $F$, and (b) $y=f(x)$ with probability $1-\eta(x)$ and $y=1-f(x)$ with probability $\eta(x)$. Here $\eta(x)$ is an unknown function that satisfies the Tsybakov noise condition with parameters $(\alpha,A)$. That is, for any $0<t\leq\frac{1}{2}$, $\eta(x)$ satisfies the condition $\Pr_{x\sim\mathcal{D}_{x}}[\eta(x)\geq\frac{1}{2}-t]\leq At^{\frac{\alpha}{1-\alpha}}$.
@@
[14, Theorem 5.1] Let $\mathcal{D}_{x}$ be a $(3,L,R,U,\beta)$-well-behaved isotropic distribution on $\mathbb{R}^{d}\times\{\pm 1\}$ that satisfies the $(\alpha,A)$-Tsybakov noise condition with respect to an unknown halfspace $f(x)=sign(\langle w,x\rangle)$. There exists an algorithm that draws $N=\beta^{4}(\frac{dUA}{RL\epsilon})^{O(1/\alpha)}log(1/\delta)$ samples from $\mathcal{D}_{x}$, runs in $poly(N,d)$ time, and computes a vector $\hat{w}$ such that, with probability $1-\delta$, we have $\Pr_{x\sim\mathcal{D}_{x}}[h(x)\neq f(x)]\leq\epsilon$.
@@
The inverse of the c.d.f. for the Bradley-Terry model $F^{-1}(x)\leq\mathrm{poly}(x)$ on $(0,\frac{1}{2}]$.
@@
The inverse of the standard logistic function $F(x)=\frac{1}{1+\exp(-x)}$ is the logit function $F^{-1}(x)=logit(x)=\ln(\frac{x}{1-x})$. Because the derivative of the logit function $\frac{1}{x-x^{2}}$ is monotonically decreasing from $\infty$ to 00 on $(0,\frac{1}{2}]$, the logit function is concave. Hence, it is bounded above by its gradient at $x=\frac{1}{2}$, which is $24x-$ 2. We have found a polynomial upper bound for $F^{-1}(x)\leq 4x-2$ for $x\in(0,\frac{1}{2}]$. ∎
@@
The inverse of the c.d.f. for the Thurstone-Mosteller model $F^{-1}(x)\leq\mathrm{poly}(x)$ on $(0,\frac{1}{2}]$.
@@ 4.2 Estimating Utility Parameters
If there is no noise, the class of linear utility functions $\mathcal{U}$ is not always passively learnable, i.e. there is a probability distribution $\mathcal{P}$ over $(x,x^{\prime})$, that no learning algorithm can achieve $e_{2}\leq\epsilon$ with probability at least $1-\delta$, for any $10<\epsilon,\delta<$ 1 under a finite set of samples.
@@
Thus, in order to achieve the second learning goal without a model assumption, we have to wish for that the incoming datapoints can help shrink the version space into a $\epsilon$-radius ball centred at $w^{*}$. The difficulty with respect to $e_{2}$ is exactly that we don’t have any control over the incoming information. Hence we will see from the following section that active learning is much more advantageous.
@@ 5.2 Estimating Utility Parameters
Input: dimension of the instance space $m$, error bound $\epsilon$
Output: $\hat{u}$ with $e_{2}(\hat{u},u)\leq\epsilon$
@@
7:     while $s>2\epsilon/\sqrt{m-1}$ do
@@
Suppose $\zeta=0$ and we can approximate the inverse of $\phi$ with $\tilde{\phi}^{-1}$ up to arbitrary precision, i.e. $||\phi(\tilde{\phi}^{-1}(x))-x||_{\inf}<c$ for any constant $c$. Then for any $\epsilon$, there is an active learning algorithm that returns a linear hypothesis $\hat{u}$ with $e_{2}(\hat{u},u)\leq\epsilon$ after asking the oracle $\mathcal{O}(m\log(\frac{\sqrt{m}}{\epsilon}))$ queries.
@@
We would like to trap the true weight $w$ inside a $m-1$ dimensional hypercube with side length $2\epsilon/\sqrt{m-1}$ so that the longest possible distance between its centre and any point in the cube is the half diagonal $\frac{\sqrt{m-1}}{2}\cdot\frac{2\epsilon}{\sqrt{m-1}}=\epsilon$. Then picking $\hat{w}$ as the centre of the cube will suffice the bounded error $e_{2}\leq\epsilon$.
@@
Depending on the noise-free signal, we can cut out at least $\frac{1}{2}-\frac{1}{10}=\frac{2}{5}$ of the space. Repeating the process for the new version space, since we go through the procedure for each dimension, we make $\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\epsilon})(m-1))$ queries in total. ∎
@@
Input: dimension of the instance space $m$, error bound $\epsilon$
Output: $\hat{u}$ with $e_{2}(\hat{u},u)\leq\epsilon$
@@
5:  $p_{0}\leftarrow F(\epsilon/\sqrt{m-1})$
@@
8:     while $d>2\epsilon/\sqrt{m-1}$ do
@@
For a fixed $\mathcal{Q}$ with corresponding c.d.f. $F$, suppose we can approximate the inverse of $\phi$ with $\tilde{\phi}^{-1}$ up to arbitrary precision, i.e. $||\phi(\tilde{\phi}^{-1}(x))-x||_{\inf}<c$ for any constant $c$, then there exists an active learning algorithm that outputs $\hat{u}$ with $e_{2}(\hat{u},u)\leq\epsilon$ with probability at least $1-\delta$ after $\mathcal{O}(\frac{1}{(p_{0}-1/2)^{2}}\log(\frac{1}{\delta}))$ queries where $p_{0}=F(\frac{\epsilon}{\sqrt{m-1}})>1/2.$
@@
Now we start calculating the upper bound of times we need to ask every turn. Let us denote the sum of a query’s labels after $T$ repetitions by $S_{T}.$ Without loss of generality, we assume that the true label of the query on $\Delta_{\phi}x$ is $1$. Then the expected value $\mathbb{E}(S_{T})>T/2$. Additionally, let $p_{0}$ denote the probability of labels flipping at the face of the hypercube to which $w$’s margin is bounded by $\frac{\epsilon}{\sqrt{m-1}}.$
@@
From Theorem 9, we know we will need $\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\epsilon})(m-1))$ accurate query labels through votes. So the cumulative confidence is to satisfy
$q^{\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\epsilon})(m-1))}\geq 1-\delta,$
where $q$ is the success rate for each query on a different hyperplane. Therefore $q\geq(1-\delta)^{\frac{1}{\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\epsilon})(m-1))}}$. And for each turn we just need $\exp(-\frac{T(p_{0}-1/2)^{2}}{4})\leq 1-q$. By a simple calculation, we get the sufficient number of repetitions from $T\geq T_{0}=-\frac{4}{(p_{0}-1/2)^{2}}\log(1-(1-\delta)^{\frac{1}{\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\epsilon})(m-1))}}).$
So our sample complexity is $\mathcal{O}(T_{0}m\log(\frac{\sqrt{m-1}}{2\epsilon})=\mathcal{O}((\frac{1}{F(\frac{\epsilon}{\sqrt{m-1}})-1/2)^{2}}\log(\frac{1}{\delta})))$. ∎
@@
For the standard logistic function $F(x)=\frac{1}{e^{-x}+1}$, the sample complexity is polynomial in $(\frac{1}{\epsilon},\log(\frac{1}{\delta}),m)$.

### v3

@@ 1 Introduction
We first investigate the “passive” Probably Approximately Correct (PAC) learning model in which pairs of inputs (outcomes) are generated according to an unknown distribution $\mathcal{P}$, and responses to these comparison queries may be corrupted by noise that comes from a given distribution $\mathcal{Q}$. We show that when there is no noise in query responses, we can indeed efficiently learn to predict responses to unseen pairwise comparison queries with respect to the true utility function in the PAC sense. However, the presence of noise with any c.d.f. continuous at 00 leads to an exponential lower bound on sample complexity in this model. Nevertheless, we show that if the input distribution is well-behaved and the query noise satisfies the Tsybakov noise condition, linear utilities are efficiently learnable in this model.
@@ 2 Related Work
Learning Utility Models from Pairwise Comparisons: The fundamental framework that we build on is random utility model learning, in which utility information is provided indirectly through ranking (e.g., pairwise) comparisons [23, 8, 35]. A common approach for learning random utility models from pairwise comparisons is by using maximum likelihood estimation (MLE) [26, 36, 35]. Unlike our setting, these typically investigate problems with a finite set of alternatives [31, 24]. Our focus, in contrast, is to learn a utility function over a continuous vector space. This problem has been considered by Zhu et al. [42], who show that linear utilities can be efficiently learned using maximum likelihood estimators in the Bradley-Terry noise model.
PAC Learning of Halfspaces With Noise: Common models of noise in random utility models have the property that the closer the pair’s difference is to the halfspace defined by the linear utility function, the higher the chance that this comparison is flipped. This kind of noise has been previously studied as boundary consistent noise in [15, 39, 38]. The possibility of an arbitrarily small difference between noise and $\frac{1}{2}$ makes learning very challenging, even for learning halfspaces [4]. Consequently, proposals for tractable learning often rely on assumptions regarding distributions of noise. The spectrum of such noise models starts from the easiest version called random classification noise (RCN) [2] – where each label is flipped independently with equal probability exactly $\eta<\frac{1}{2}$ – where efficient algorithms exist [6]. At the other end are malicious noise [19] and agnostic learning [20, 21, 11, 13], where an adversarial oracle can flip any small constant fraction of the labels, typically making these models less tractable. Additionally, active learning has been proposed as a solution to reduce the number of samples the learner needs to query before approximately learning the concept class with high probability. While active learning requires exponentially fewer labeled samples than PAC-learning for simple classes such as thresholds in one dimension, it fails in general to provide asymptotic improvement for broad classes such as halfspaces [12].
@@ 3 Preliminaries
We further define the difference between a pair $(\phi(x),\phi(x^{\prime}))$ as $\Delta_{\phi}(x)\equiv(\phi(x^{\prime})-\phi(x))\in[-1,1]^{m}$. Since $\tilde{u}(x^{\prime})-\tilde{u}(x)=w^{T}\Delta_{\phi}(x)+\tilde{\zeta}(\phi(x^{\prime}))-\tilde{\zeta}(\phi(x))$, it will be most useful to consider $\zeta\equiv\tilde{\zeta}(\phi(x^{\prime}))-\tilde{\zeta}(\phi(x))$ for a pair of feature vectors $(\phi(x),\phi(x^{\prime}))$, so that $Z_{\tilde{u}}(x,x^{\prime})=\mathrm{sign}(w^{T}\Delta_{\phi}(x)+\zeta)$. Let $\mathcal{Q}$ be the distribution over $\zeta$ induced by $\tilde{\mathcal{Q}}$, and let $F(\zeta)$ be its c.d.f., which we assume is continuous on $\mathbb{R}$ except in the noise-free special case. Then the probability that $y=1$ for a pair $(\phi(x),\phi(x^{\prime}))$, that is, the probability that $x^{\prime}\succ x$, is $\Pr(x^{\prime}\succ x)=\Pr(\zeta\leq w^{T}\Delta_{\phi}(x))=F(w^{T}\Delta_{\phi}(x))$. As we assume $\Pr(x^{\prime}\succ x)+\Pr(x\succ x^{\prime})=1$, $F(w^{T}\Delta_{\phi}(x))+F(-w^{T}\Delta_{\phi}(x))=1$ leading to $F(x)+F(-x)=1$ and $F(0)=1/2$.
@@
Our learnability discussion will be based on an adaptation of the Probably Approximately Correct (PAC) Learning framework [33], with the goal of identifying a utility model $\hat{u}$ that has error at most $\varepsilon$ with probability at least $1-\delta$ for $\varepsilon>0$ and $\delta>0$. We now formalize this as PAC learnability from pairwise comparisons (PAC-PC). Let $\mathbb{N}$ be the space of natural numbers.
Definition 1 (PAC-PC learnability).
Given a noise distribution $\mathcal{Q}$, a utility class $\mathcal{U}$ is PAC learnable from pairwise comparisons (PAC-PC learnable) for error function $e(\hat{u},u)$ if there is a learning algorithm $\mathcal{A}$ and a function $n_{A}:(0,1)^{2}\rightarrow\mathbb{N}$ such that for any input distribution $\mathcal{P}$, and $\forall\varepsilon,\delta\in(0,1)$, whenever $n\geq n_{A}(\varepsilon,\delta)$, $\mathcal{A}(\{(x_{i},x_{i}^{\prime},y_{i})\}_{i=1}^{n})$ with $(x_{i},x_{i}^{\prime})\sim\mathcal{P}$ i.i.d. and $y_{i}=\mathrm{sign}(u(x_{i}^{\prime})-u(x_{i})+\zeta_{i})$ for $\zeta_{i}\sim\mathcal{Q}$ i.i.d. returns $\hat{u}$ with probability at least $1-\delta$ such that $e(\hat{u},u)\leq\varepsilon$. Moreover, if $n_{A}$ is polynomial in $1/\varepsilon,1/\delta,\mathrm{Dim}(\mathcal{U})$, where $\mathrm{Dim}(\mathcal{U})$ is the dimension of $\mathcal{U}$, then we say $\mathcal{U}$ is efficiently PAC-PC learnable.
@@ 4.1 Predicting Pairwise Preferences
Suppose $\zeta=0$. Then linear utility functions are efficiently PAC-PC learnable under the error function $e_{1}$.
@@
Let $y=g(\Delta_{\phi}(x))=\mathrm{sign}(w^{*^{T}}\Delta_{\phi}(x))$. Then PAC-PC learning under $e_{1}$ and $\mathcal{P}$ is equivalent to PAC learnability of halfspaces $g(\Delta_{\phi}(x))$ under $\mathcal{P}^{\prime}_{\phi}$. As shown by Blumer et al. [7], halfspaces are learnable with sample complexity $\mathcal{O}(\frac{1}{\varepsilon}(m+\log(\frac{1}{\delta})))$. ∎
@@
Due to the continuity of $F$ at 00, for any $t>0$, there exists $s>0$ such that for all $|v-0|\leq s$, $|F(v)-F(0)|\leq t$. As $F$ is monotonically increasing, and $-|w^{*^{T}}\Delta_{\phi}(x)|<0$, $\eta(\Delta_{\phi}(x))<F(0)=\frac{1}{2}$. Hence for a fixed $\varepsilon<1$, we can pick $t_{0}=1/\exp(1/\varepsilon)$. Then there exists a corresponding $s_{0}$ such that for a $(\phi(x),\phi(x^{\prime}))$ with $\Delta_{\phi}x$ of norm at most $s_{0}$, for all $w$, $|w^{T}\Delta_{\phi}(x)-0|\leq s_{0}$, $|\eta(\Delta_{\phi}(x))-\frac{1}{2}|\leq t_{0}.$ Now consider the oracle who always chooses one such pair $(\phi(x),\phi(x^{\prime}))$. With fixed probability at most $\frac{1}{2}+t_{0}$ it outputs the true label and with probability $\frac{1}{2}-t_{0}$ it outputs the false label. In order to establish the Binomial distribution’s confidence over $1-\delta$, we need sample complexity $\tilde{\Omega}(\frac{1}{t_{0}})=\tilde{\Omega}(\exp(\frac{1}{\varepsilon}))$, which is exponential in $\frac{1}{\varepsilon}$.∎
@@
[14] Let $C$ be a concept class of Boolean-valued functions over $X=\mathbb{R}^{d}$, $F$ be a family of distributions on $X$, $10<\varepsilon<$ 1 be the error parameter, and $10\leq\alpha<$ 1, $A>0$ be parameters of the noise model. Let $f$ be an unknown target function in $C$. A Tsybakov example oracle, $EX^{Tsyb}(f,F)$, works as follows: Each time $EX^{Tsyb}(f,F)$ is invoked, it returns a labeled example $(x,y)$, such that: (a) $x\sim\mathcal{D}_{x}$, where $\mathcal{D}_{x}$ is a fixed distribution in $F$, and (b) $y=f(x)$ with probability $1-\eta(x)$ and $y=1-f(x)$ with probability $\eta(x)$. Here $\eta(x)$ is an unknown function that satisfies the Tsybakov noise condition with parameters $(\alpha,A)$. That is, for any $0<t\leq\frac{1}{2}$, $\eta(x)$ satisfies the condition $\Pr_{x\sim\mathcal{D}_{x}}[\eta(x)\geq\frac{1}{2}-t]\leq At^{\frac{\alpha}{1-\alpha}}$.
@@
[14, Theorem 5.1] Let $\mathcal{D}_{x}$ be a $(3,L,R,U,\beta)$-well-behaved isotropic distribution on $\mathbb{R}^{d}\times\{\pm 1\}$ that satisfies the $(\alpha,A)$-Tsybakov noise condition with respect to an unknown halfspace $f(x)=sign(\langle w,x\rangle)$. There exists an algorithm that draws $N=\beta^{4}(\frac{dUA}{RL\varepsilon})^{O(1/\alpha)}log(1/\delta)$ samples from $\mathcal{D}_{x}$, runs in $poly(N,d)$ time, and computes a vector $\hat{w}$ such that, with probability $1-\delta$, we have $\Pr_{x\sim\mathcal{D}_{x}}[h(x)\neq f(x)]\leq\varepsilon$.
@@
The inverse of the c.d.f. for the Bradley-Terry model satisfies $F^{-1}(x)\leq\mathrm{poly}(x)$ on $(0,\frac{1}{2}]$.
@@
The inverse of the standard logistic function $F(x)=\frac{1}{1+\exp(-x)}$ is the logit function $F^{-1}(x)=logit(x)=\ln(\frac{x}{1-x})$. Because the derivative of the logit function $\frac{1}{x-x^{2}}$ is monotonically decreasing from $\infty$ to $4$ on $(0,\frac{1}{2}]$, the logit function is concave. Hence, it is bounded above by its gradient at $x=\frac{1}{2}$, which is $24x-$ 2. We have found a polynomial upper bound for $F^{-1}(x)\leq 4x-2$ for $x\in(0,\frac{1}{2}]$. ∎
@@
The inverse of the c.d.f. for the Thurstone-Mosteller model satisfies $F^{-1}(x)\leq\mathrm{poly}(x)$ on $(0,\frac{1}{2}]$.
@@ 4.2 Estimating Utility Parameters
If there is no noise, the class of linear utility functions $\mathcal{U}$ is not always passively learnable, i.e. there is a probability distribution $\mathcal{P}$ over $(x,x^{\prime})$, that no learning algorithm can achieve $e_{2}\leq\varepsilon$ with probability at least $1-\delta$, for any $10<\varepsilon,\delta<$ 1 under a finite set of samples.
@@
Thus, in order to achieve the second learning goal without a model assumption, we have to wish for that the incoming datapoints can help shrink the version space into a $\varepsilon$-radius ball centred at $w^{*}$. The difficulty with respect to $e_{2}$ is exactly that we don’t have any control over the incoming information. Hence we will see from the following section that active learning is much more advantageous.
@@ 5.2 Estimating Utility Parameters
Input: dimension of the instance space $m$, error bound $\varepsilon$
Output: $\hat{u}$ with $e_{2}(\hat{u},u)\leq\varepsilon$
@@
7:     while $s>2\varepsilon/\sqrt{m-1}$ do
@@
Suppose $\zeta=0$ and we can approximate the inverse of $\phi$ with $\tilde{\phi}^{-1}$ up to arbitrary precision, i.e. $||\phi(\tilde{\phi}^{-1}(x))-x||_{\inf}<c$ for any constant $c$. Then for any $\varepsilon$, there is an active learning algorithm that returns a linear hypothesis $\hat{u}$ with $e_{2}(\hat{u},u)\leq\varepsilon$ after asking the oracle $\mathcal{O}(m\log(\frac{\sqrt{m}}{\varepsilon}))$ queries.
@@
We would like to trap the true weight $w$ inside a $m-1$ dimensional hypercube with side length $2\varepsilon/\sqrt{m-1}$ so that the longest possible distance between its centre and any point in the cube is the half diagonal $\frac{\sqrt{m-1}}{2}\cdot\frac{2\varepsilon}{\sqrt{m-1}}=\varepsilon$. Then picking $\hat{w}$ as the centre of the cube will suffice the bounded error $e_{2}\leq\varepsilon$.
@@
Depending on the noise-free signal, we can cut out at least $\frac{1}{2}-\frac{1}{10}=\frac{2}{5}$ of the space. Repeating the process for the new version space, since we go through the procedure for each dimension, we make $\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\varepsilon})(m-1))$ queries in total. ∎
@@
Input: dimension of the instance space $m$, error bound $\varepsilon$
Output: $\hat{u}$ with $e_{2}(\hat{u},u)\leq\varepsilon$
@@
5:  $p_{0}\leftarrow F(\varepsilon/\sqrt{m-1})$
@@
8:     while $d>2\varepsilon/\sqrt{m-1}$ do
@@
For a fixed $\mathcal{Q}$ with corresponding c.d.f. $F$, suppose we can approximate the inverse of $\phi$ with $\tilde{\phi}^{-1}$ up to arbitrary precision, i.e. $||\phi(\tilde{\phi}^{-1}(x))-x||_{\inf}<c$ for any constant $c$, then there exists an active learning algorithm that outputs $\hat{u}$ with $e_{2}(\hat{u},u)\leq\varepsilon$ with probability at least $1-\delta$ after $\mathcal{O}(\frac{1}{(p_{0}-1/2)^{2}}\log(\frac{1}{\delta}))$ queries where $p_{0}=F(\frac{\varepsilon}{\sqrt{m-1}})>1/2.$
@@
Now we start calculating the upper bound of times we need to ask every turn. Let us denote the sum of a query’s labels after $T$ repetitions by $S_{T}.$ Without loss of generality, we assume that the true label of the query on $\Delta_{\phi}x$ is $1$. Then the expected value $\mathbb{E}(S_{T})>T/2$. Additionally, let $p_{0}$ denote the probability of labels flipping at the face of the hypercube to which $w$’s margin is bounded by $\frac{\varepsilon}{\sqrt{m-1}}.$
@@
From Theorem 9, we know we will need $\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\varepsilon})(m-1))$ accurate query labels through votes. So the cumulative confidence is to satisfy
$q^{\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\varepsilon})(m-1))}\geq 1-\delta,$
where $q$ is the success rate for each query on a different hyperplane. Therefore $q\geq(1-\delta)^{\frac{1}{\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\varepsilon})(m-1))}}$. And for each turn we just need $\exp(-\frac{T(p_{0}-1/2)^{2}}{4})\leq 1-q$. By a simple calculation, we get the sufficient number of repetitions from $T\geq T_{0}=-\frac{4}{(p_{0}-1/2)^{2}}\log(1-(1-\delta)^{\frac{1}{\mathcal{O}(\log(\frac{\sqrt{m-1}}{2\varepsilon})(m-1))}}).$
So our sample complexity is $\mathcal{O}(T_{0}m\log(\frac{\sqrt{m-1}}{2\varepsilon})=\mathcal{O}((\frac{1}{F(\frac{\varepsilon}{\sqrt{m-1}})-1/2)^{2}}\log(\frac{1}{\delta})))$. ∎
@@
For the standard logistic function $F(x)=\frac{1}{e^{-x}+1}$, the sample complexity is polynomial in $(\frac{1}{\varepsilon},\log(\frac{1}{\delta}),m)$.